OpenStudy (anonymous):

Which is the solution to 3x – 2(x + 2) = –4x – 4 + 5x? A. –4 B. no solution C. 2 D. all real numbers

OpenStudy (anonymous):

can anyone please help?

OpenStudy (anonymous):

3x - 2x -4 = x-4

OpenStudy (anonymous):

is this correct?

OpenStudy (anonymous):

x-4 = x-4 right?

OpenStudy (andriod09):

Yes, that's correct RH.

OpenStudy (anonymous):

so which of the answers would be correct?

OpenStudy (andriod09):

Np @SharpEdge Also, RH, after you get x-4=x-4, what do you do after that? That's not the answer, just so you're aware. Because you're trying to simplify the term "x" or to get it to stand along, thus, having more than one "x" in the equation means its not done.

OpenStudy (anonymous):

no solution?

OpenStudy (andriod09):

That's correct, if you have an equation where you already tell that you will have no "x" when you have to find "x" like this one, which is the one you had: \[3x – 2(x + 2) = –4x – 4 + 5x\], you can easily see that it wouldn't really work out. Because just looking at it, if you distribute in your head, you would realize that \[3x-2x+2=1x+4\] thus making it to where you wouldn't be able to simplify the answer past 1x.

OpenStudy (anonymous):

Thanks!!!!

OpenStudy (andriod09):

No problem. Any time.